OpenStudy (anonymous):

How do I find the simplest radical form of 75^1/2 divided by 3^1/2+ 8^1/2?

OpenStudy (anonymous):

hey how do u do the pink bar i know there is some code

OpenStudy (anonymous):

unfortunately , I am stuck with this based on my level

OpenStudy (anonymous):

kkk

OpenStudy (anonymous):

\[\frac{\sqrt{75}}{\sqrt{8} +\sqrt{3}}\] \[\frac{\sqrt{25}\sqrt{3}}{\sqrt{8} +\sqrt{3}}\] \[\frac{5\sqrt{3}}{\sqrt{8} +\sqrt{3}}\]

OpenStudy (anonymous):

hopefully , I didn't screw up again

OpenStudy (anonymous):

Thanks :)

OpenStudy (anonymous):

now multiply the numerator and denominator by the conjugate\[\sqrt{8}-\sqrt{3}\]

OpenStudy (anonymous):

\[5\sqrt{24}-15 / 5\]

OpenStudy (anonymous):

\[2\sqrt{6}-3\]